# Verifiable Fine-Tuning for LLMs: Zero-Knowledge Training Proofs Bound to Data Provenance and Policy

### v2

@@ II-J Summary of Gaps
Despite vibrant progress, several gaps persist. Proof construction for training remains expensive at LLM scale; while arithmetization strategies and series expansions are improving [sun2023zkdl, taylorZKML2025], end-to-end costs will likely require recursive aggregation and amortized audit strategies as proposed in decentralized contexts [probAudit2025]. Provenance capture, even when automated, is only as good as what pipelines expose; integrating commit-and-sample protocols at the data interface remains an open systems challenge [schlegel2023mlflow2prov]. Watermarks can provide useful downstream signals but cannot shoulder compliance alone in adversarial settings [jovanovic2024wmstealing, zhang2024mipSteal]. And unlearning with verifiability is still embryonic, requiring updateable proofs and revocation semantics to maintain coherent audit trails [eisenhofer2022unlearning]. Our framework addresses these gaps by co-designing commitments at the data boundary, parameter-efficient proof circuits for updates, and aggregation-friendly verification tailored to federated and edge deployments, while exposing interfaces that harmonize with property-card attestations from TEEs [duddu2025laminator] and with existing provenance infrastructures [schlegel2025isProvenance].
In summary, prior work provides strong building blocks—PoL-style pre-commitment [jia2021proof, zhao2024polIncentive], ZK-ML circuits for training and LLMs [abbaszadeh2024zkpot, sun2024zkllm], provenance capture and distributional attestation [schelter2024provenanceScreen, dold2024attesting], and federated systems that respect heterogeneity and resource limits [yao2024ferrari, liu2021communication]. We synthesize these lines into a coherent VFT protocol aimed at making fine-tuning demonstrably correct, private, and auditable in the settings where it matters most.
@@ III-C Formal Statement of the Verified Claim
$\theta_{t+1}$ $=\theta_{t}-\eta_{t}\,\frac{\hat{m}_{t+1}}{\sqrt{\hat{v}_{t+1}}+\epsilon}-\eta_{t}\lambda\theta_{t}\,$,
@@ III-D Dataset Commitment and Admissible Sampling
$\sum_{t\in\text{epoch}}Me_{t}\;\preceq\;\gamma\,$, (3)
@@ III-F Quantization and Fixed-Field Representation
$\bar{x}\;=\;\left\lfloor Sx\right\rceil,\quad\overline{xy}\approx\frac{\bar{x}\cdot\bar{y}}{S}\,$, (5)
@@ IV-A Datasets
Instruction-following corpora.
For text-only fine-tuning, we construct two instruction mixtures. Mix-A targets English-centric instruction following with a spectrum of reasoning, summarization, and tool-use exemplars. Mix-B is bilingual (English–Chinese) and includes classification, extraction, and generation tasks to probe cross-lingual behavior. Both mixtures adopt a consistent conversational schema (system, user, assistant) and normalize whitespace, Unicode punctuation, and code blocks to stabilize tokenization. We cap context length at 4,096 tokens unless otherwise stated. To emulate realistic governance, Mix-B includes a small fraction of domain-tagged items (e.g., medical, finance, law) used purely for bin-level counters in policy checks; raw content is not disclosed under private-sampling mode.
@@
Unless stated otherwise: Mix-A contains 200,000 instruction–response pairs, Mix-B 120,000 pairs (60% English, 40% Chinese), safety slices add 15,000 pairs, and synthetic telemetry descriptors add 5,000 items. Sequences are right-padded to 2,048 tokens for PEFT runs targeting lower proof cost; long-context ablations use 4,096 tokens. We report how proof cost scales with sequence length.
@@ IV-C1 Model Quality
We report validation perplexity (PPL) on a structured held-out corpus and Expected Calibration Error (ECE) computed over next-token probabilities on random sub-samples. Since VFT introduces fixed-point approximations and polynomial surrogates, we track the shift in PPL relative to a non-proved PEFT baseline under the same hyperparameters. Our acceptance target is a PPL increase no greater than $0.5$ absolute or $3\%$ relative for $r\in\{8,16\}$ at 2,0482{,}048 tokens; long-context runs tolerate a slightly larger budget due to increased approximation pressure.
@@ IV-D Ablation and Sensitivity Protocols
We repeat selected runs with 4,096-token sequences to quantify scaling with context length. Because attention matmuls scale quadratically in sequence length per head and our circuits include row-sum checks and lookup normalizations, we expect superlinear growth in constraints; we report the exact slope and whether batching strategies mitigate the cost.
@@ IV-F Acceptance Targets
  • Model quality: For $r\in\{8,16\}$ at 2,048 tokens, VFT runs should maintain instruction-following metrics within 0.3 absolute points of a non-proved PEFT baseline and increase PPL by at most 3% relative.
@@ V Results & Discussion
We evaluate Verifiable Fine-Tuning (VFT) along three axes: (i) utility—instruction-following quality, perplexity, calibration, and safety; (ii) verification performance—prover time, verifier time, proof size, and aggregation efficiency; and (iii) policy & privacy—quota compliance and leakage under private sampling. Results are reported for the standard configuration described in Section Experimental Setup and are organized as main outcomes, ablations, and visual analyses. Where helpful, we contextualize findings with prior work on proof-of-learning, verifiable ML, provenance, watermarking, and federated learning [jia2021proof, sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm, schelter2024provenanceScreen, schlegel2025isProvenance, kirchenbauer2023watermark, liu2024textwmSurvey, liu2021communication, yao2024ferrari]. Unless otherwise noted, all reported values are the average of three seeds with standard deviations; detailed per-task breakdowns appear in the appendix. We refer to the system overview in Fig. 1 and summarize key outcomes in Table I, ablations in Table III, and verifier costs in Table II.
TABLE I: Main results on Mix-A / Mix-B (LoRA rank $r\in\{8,16\}$, seq=2,048).
@@ V-A Main Utility Outcomes
On Mix-A, VFT with LoRA rank r=8r{=}8 and sequence length 2,048 achieves near-parity with the non-proved PEFT baseline: ROUGE-L decreases by $0.12\pm 0.05$ points on summarization prompts, while extraction BLEU drops by $0.09\pm 0.04$. Open-ended win-rate against the baseline evaluator is $49.6\%\pm 0.8$ (vs. the inherent $50\%$ tie target), suggesting the fixed-point and approximation budgets do not materially harm generation quality. On Mix-B, bilingual performance shows an English–Chinese delta of 0.30.30.50.5 points on task-weighted averages, consistent with cross-lingual PEFT behavior and within our predefined acceptance budget. These results contrast with early ZK-ML for inference only, where polynomial approximations sometimes degraded accuracy substantially [liu2021zkcnn, wang2023ezdps]; by targeting PEFT subspaces and adopting tight error budgets with lookup-accelerated non-linearities, VFT maintains utility while generating proofs for training.
@@ V-B Verification Performance
Per-step prover time on Tier-P2 with r=8r{=}8, 2,048 tokens averages $16.8\pm 0.9$ s; for r=16r{=}16 it is $31.2\pm 1.4$ s. Peak memory is $23.6\pm 1.1$ GB and $37.9\pm 1.5$ GB, respectively. Constraint counts scale near-linearly with $P_{\mathrm{PEFT}}$ and superlinearly with sequence length due to attention checks; the slope is lower than naive quadratic because our lookup-based softmax imposes row-sum constraints rather than polynomial expansions per entry. Fig. 4 (referenced) shows a breakdown: witness generation consumes $\sim$35% of time, polynomial/FFT evaluation $\sim$28%, MSM/FRI $\sim$30%, and overheads the remainder. Compared to training-proof prototypes that circuitize dense updates [sun2023zkdl, abbaszadeh2024zkpot], our PEFT restriction reduces constraint volume by an order of magnitude for typical transformer blocks.
@@
Folding T=10, [^0] T{=}10{,}000 step proofs into a single epoch certificate with 4 GPUs takes $8.7\pm 0.4$ minutes; doubling GPUs yields sub-linear speed-ups due to memory bandwidth contention in MSM/FFT kernels. Fig. 5 (referenced) plots depth vs. wall-clock; beyond 4 GPUs, I/O streaming of witnesses becomes the bottleneck, motivating future work on incremental transcripts and pipeline-parallel provers.
@@ V-C Policy and Privacy Outcomes
TABLE III: Ablations on Mix-A (default: r=8r{=}8, q=14q{=}14, $\delta_{\mathrm{sm}}{=}3{\times}10^{-3}$, seq=2,048=2{,}048, public sampling). $\downarrow$ lower is better, $\uparrow$ higher is better.
@@ V-D Ablations
Increasing LoRA rank from r=4r{=}4 to r=16r{=}16 improves instruction metrics by $\sim$0.6 points but doubles prover cost; placing adapters in attention and MLP layers yields modest quality gains (+0.2+0.20.30.3 points) for $\sim$25% higher cost. A normalized metric—proof-time per trainable parameter—is nearly constant across ranks, indicating linearity in $P_{\mathrm{PEFT}}$. This supports our design choice of PEFT-restricted circuits and resonates with recent practice that parameter-efficient fine-tuning dominates real deployments [10251628, yao2024ferrari, yan2024peaches].
@@
Moving from 2,048 to 4,096 tokens increases constraint counts by $2.23\times$ and prover time by $2.38\times$ at fixed rank; PPL improves slightly due to longer contexts. Attention matmuls drive the increase; lookup-based softmax keeps the growth sub-quadratic relative to naive polynomialization. Fig. 7 (referenced) shows log–log slopes and confidence intervals. These costs outpace those reported for inference-only ZK [wang2023ezdps] because training proofs include backward passes and optimizer updates.
@@ V-E Comparisons to Prior Art
Watermarks are complementary post-release signals [kirchenbauer2023watermark, wen2023treering]. However, their brittleness under adaptive attacks [gu2024learnability, jovanovic2024wmstealing, zhang2024mipSteal] and mixed success in practice [dathathri2024natureWatermark, liu2024textwmSurvey] make them insufficient for compliance. Our manifest can record where (if at all) watermarks are inserted, aiding downstream detectors; but the training story rests on proofs, not beacons. Robust watermark designs [zhao2023provable, qu2025multibit] help on the output side and can be paired with VFT for defense-in-depth.
@@
Decentralized and edge learning has seen sustained progress on communication and personalization [liu2021communication, liu2023yoga, yao2024ferrari, yan2024peaches, 10508334, 10251628, liu2024enhancing]. VFT complements these lines with verifiability: the aggregator can accept only updates accompanied by partial proofs or coverage certificates; clients can hide indices in private-sampling mode. Our experiments demonstrate feasibility within the bandwidth and latency envelopes typical of enterprise deployments.
@@ V-K Practical Guidance
If proof budgets are tight, r=8r{=}8 delivers a favorable utility–cost balance for 2,048-token contexts. For high-stakes domains requiring long contexts, invest in r=16r{=}16 but limit softmax tolerances to keep fidelity.
@@ V-L Summary of Findings
VFT delivers cryptographically strong, privacy-preserving assurances about the fine-tuning trajectory while preserving model utility within tight budgets. Compared to the state of the art in proof-of-training and zk-LLM [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm], our PEFT-centric circuits and approximation budgeting reduce prover cost and maintain accuracy. By binding provenance manifests to computation-level proofs, VFT operationalizes calls for accountable data governance [longpre2024data, schelter2024provenanceScreen, schlegel2025isProvenance] and sidesteps the brittleness of watermark-only accountability [kirchenbauer2023watermark, liu2024textwmSurvey]. Federated evaluations confirm compatibility with heterogeneous devices and probabilistic audits [liu2021communication, yao2024ferrari, probAudit2025]. Across datasets and settings, quota compliance is perfect, index leakage is negligible, and verification costs fit within the latency envelopes of practical audits. These results support the view that verifiable fine-tuning can be deployed today in centralized and decentralized pipelines to close the trust gap around LLM adaptation.

### v3

@@ II-J Summary of Gaps
Despite vibrant progress, several gaps persist. Proof construction for training remains expensive at LLM scale; while arithmetization strategies and series expansions are improving [sun2023zkdl, taylorZKML2025], end-to-end costs will likely require recursive aggregation and amortized audit strategies as proposed in decentralized contexts [probAudit2025]. Provenance capture, even when automated, is only as good as what pipelines expose; integrating commit-and-sample protocols at the data interface remains an open systems challenge [schlegel2023mlflow2prov]. Watermarks can provide useful downstream signals but cannot shoulder compliance alone in adversarial settings [jovanovic2024wmstealing, zhang2024mipSteal]. And unlearning with verifiability is still embryonic, requiring updateable proofs and revocation semantics to maintain coherent audit trails [eisenhofer2022unlearning]. Our framework addresses these gaps by co-designing commitments at the data boundary, parameter-efficient proof circuits for updates, and aggregation-friendly verification tailored to federated and edge deployments, while exposing interfaces that harmonize with property-card attestations from TEEs and with existing provenance infrastructures [schlegel2025isProvenance].
In summary, prior work provides strong building blocks—PoL-style pre-commitment [jia2021proof, zhao2024polIncentive], ZK-ML circuits for training and LLMs [abbaszadeh2024zkpot, sun2024zkllm], provenance capture and distributional attestation [schelter2024provenanceScreen, dold2024attesting], and federated systems that respect heterogeneity and resource limits . We synthesize these lines into a coherent VFT protocol aimed at making fine-tuning demonstrably correct, private, and auditable in the settings where it matters most.
@@ III-C Formal Statement of the Verified Claim
$\theta_{t+1}$ $=\theta_{t}-\eta_{t}\,\frac{\hat{m}_{t+1}}{\sqrt{\hat{v}_{t+1}}+\epsilon}-\eta_{t}\lambda\theta_{t}\,$
@@ III-D Dataset Commitment and Admissible Sampling
$\sum_{t\in\text{epoch}}Me_{t}\;\preceq\;\gamma\,$ (3)
@@ III-F Quantization and Fixed-Field Representation
$\bar{x}\;=\;\left\lfloor Sx\right\rceil,\quad\overline{xy}\approx\frac{\bar{x}\cdot\bar{y}}{S}\,$ (5)
@@ IV-A Datasets
Unless stated otherwise: Mix-A contains 200000 instruction–response pairs, Mix-B 120000 pairs (60% English, 40% Chinese), safety slices add 15000 pairs, and synthetic telemetry descriptors add 5000 items. Sequences are right-padded to 2048 tokens for PEFT runs targeting lower proof cost; long-context ablations use 4096 tokens. We report how proof cost scales with sequence length.
@@ IV-C1 Model Quality
We report validation perplexity (PPL) on a structured held-out corpus and Expected Calibration Error (ECE) computed over next-token probabilities on random sub-samples. Since VFT introduces fixed-point approximations and polynomial surrogates, we track the shift in PPL relative to a non-proved PEFT baseline under the same hyperparameters. Our acceptance target is a PPL increase no greater than $0.5$ absolute or $3\%$ relative for $r\in\{8,16\}$ at $2048$ tokens; long-context runs tolerate a slightly larger budget due to increased approximation pressure.
@@ IV-D Ablation and Sensitivity Protocols
We repeat selected runs with 4096-token sequences to quantify scaling with context length. Because attention matmuls scale quadratically in sequence length per head and our circuits include row-sum checks and lookup normalizations, we expect superlinear growth in constraints; we report the exact slope and whether batching strategies mitigate the cost.
@@ IV-F Acceptance Targets
  • Model quality: For $r\in\{8,16\}$ at 2048 tokens, VFT runs should maintain instruction-following metrics within 0.3 absolute points of a non-proved PEFT baseline and increase PPL by at most 3% relative.
@@ V Results & Discussion
We evaluate Verifiable Fine-Tuning (VFT) along three axes: (i) utility—instruction-following quality, perplexity, calibration, and safety; (ii) verification performance—prover time, verifier time, proof size, and aggregation efficiency; and (iii) policy & privacy—quota compliance and leakage under private sampling. Results are reported for the standard configuration described in Section Experimental Setup and are organized as main outcomes, ablations, and visual analyses. Where helpful, we contextualize findings with prior work on proof-of-learning, verifiable ML, provenance, watermarking, and federated learning [jia2021proof, sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm, schelter2024provenanceScreen, schlegel2025isProvenance, kirchenbauer2023watermark, liu2024textwmSurvey]. Unless otherwise noted, all reported values are the average of three seeds with standard deviations; detailed per-task breakdowns appear in the appendix. We refer to the system overview in Fig. 1 and summarize key outcomes in Table I, ablations in Table III, and verifier costs in Table II.
TABLE I: Main results on Mix-A / Mix-B (LoRA rank $r\in\{8,16\}$, seq=2048).
@@ V-A Main Utility Outcomes
On Mix-A, VFT with LoRA rank r=8r{=}8 and sequence length 2048 achieves near-parity with the non-proved PEFT baseline: ROUGE-L decreases by $0.12\pm 0.05$ points on summarization prompts, while extraction BLEU drops by $0.09\pm 0.04$. Open-ended win-rate against the baseline evaluator is $49.6\%\pm 0.8$ (vs. the inherent $50\%$ tie target), suggesting the fixed-point and approximation budgets do not materially harm generation quality. On Mix-B, bilingual performance shows an English–Chinese delta of 0.30.30.50.5 points on task-weighted averages, consistent with cross-lingual PEFT behavior and within our predefined acceptance budget. These results contrast with early ZK-ML for inference only, where polynomial approximations sometimes degraded accuracy substantially [liu2021zkcnn, wang2023ezdps]; by targeting PEFT subspaces and adopting tight error budgets with lookup-accelerated non-linearities, VFT maintains utility while generating proofs for training.
@@ V-B Verification Performance
Per-step prover time on Tier-P2 with r=8r{=}8, 2048 tokens averages $16.8\pm 0.9$ s; for r=16r{=}16 it is $31.2\pm 1.4$ s. Peak memory is $23.6\pm 1.1$ GB and $37.9\pm 1.5$ GB, respectively. Constraint counts scale near-linearly with $P_{\mathrm{PEFT}}$ and superlinearly with sequence length due to attention checks; the slope is lower than naive quadratic because our lookup-based softmax imposes row-sum constraints rather than polynomial expansions per entry. Fig. 4 (referenced) shows a breakdown: witness generation consumes $\sim$35% of time, polynomial/FFT evaluation $\sim$28%, MSM/FRI $\sim$30%, and overheads the remainder. Compared to training-proof prototypes that circuitize dense updates [sun2023zkdl, abbaszadeh2024zkpot], our PEFT restriction reduces constraint volume by an order of magnitude for typical transformer blocks.
@@
Folding T=10000T{=}10000 step proofs into a single epoch certificate with 4 GPUs takes $8.7\pm 0.4$ minutes; doubling GPUs yields sub-linear speed-ups due to memory bandwidth contention in MSM/FFT kernels. Fig. 5 (referenced) plots depth vs. wall-clock; beyond 4 GPUs, I/O streaming of witnesses becomes the bottleneck, motivating future work on incremental transcripts and pipeline-parallel provers.
@@ V-C Policy and Privacy Outcomes
TABLE III: Ablations on Mix-A (default: r=8r{=}8, q=14q{=}14, $\delta_{\mathrm{sm}}{=}3{\times}10^{-3}$, seq=2048=2048, public sampling). $\downarrow$ lower is better, $\uparrow$ higher is better.
@@ V-D Ablations
Increasing LoRA rank from r=4r{=}4 to r=16r{=}16 improves instruction metrics by $\sim$0.6 points but doubles prover cost; placing adapters in attention and MLP layers yields modest quality gains (+0.2+0.20.30.3 points) for $\sim$25% higher cost. A normalized metric—proof-time per trainable parameter—is nearly constant across ranks, indicating linearity in $P_{\mathrm{PEFT}}$. This supports our design choice of PEFT-restricted circuits and resonates with recent practice that parameter-efficient fine-tuning dominates real deployments .
@@
Moving from 2048 to 4096 tokens increases constraint counts by $2.23\times$ and prover time by $2.38\times$ at fixed rank; PPL improves slightly due to longer contexts. Attention matmuls drive the increase; lookup-based softmax keeps the growth sub-quadratic relative to naive polynomialization. Fig. 7 (referenced) shows log–log slopes and confidence intervals. These costs outpace those reported for inference-only ZK [wang2023ezdps] because training proofs include backward passes and optimizer updates.
@@ V-E Comparisons to Prior Art
Watermarks are complementary post-release signals [kirchenbauer2023watermark, wen2023treering]. However, their brittleness under adaptive attacks [gu2024learnability, jovanovic2024wmstealing, zhang2024mipSteal] and mixed success in practice [dathathri2024natureWatermark, liu2024textwmSurvey] make them insufficient for compliance. Our manifest can record where (if at all) watermarks are inserted, aiding downstream detectors; but the training story rests on proofs, not beacons. Robust watermark designs [qu2025multibit] help on the output side and can be paired with VFT for defense-in-depth.
@@
Decentralized and edge learning has seen sustained progress on communication and personalization . VFT complements these lines with verifiability: the aggregator can accept only updates accompanied by partial proofs or coverage certificates; clients can hide indices in private-sampling mode. Our experiments demonstrate feasibility within the bandwidth and latency envelopes typical of enterprise deployments.
@@ V-K Practical Guidance
If proof budgets are tight, r=8r{=}8 delivers a favorable utility–cost balance for 2048-token contexts. For high-stakes domains requiring long contexts, invest in r=16r{=}16 but limit softmax tolerances to keep fidelity.
@@ V-L Summary of Findings
VFT delivers cryptographically strong, privacy-preserving assurances about the fine-tuning trajectory while preserving model utility within tight budgets. Compared to the state of the art in proof-of-training and zk-LLM [sun2023zkdl, abbaszadeh2024zkpot, sun2024zkllm], our PEFT-centric circuits and approximation budgeting reduce prover cost and maintain accuracy. By binding provenance manifests to computation-level proofs, VFT operationalizes calls for accountable data governance [longpre2024data, schelter2024provenanceScreen, schlegel2025isProvenance] and sidesteps the brittleness of watermark-only accountability [kirchenbauer2023watermark, liu2024textwmSurvey]. Federated evaluations confirm compatibility with heterogeneous devices and probabilistic audits [probAudit2025]. Across datasets and settings, quota compliance is perfect, index leakage is negligible, and verification costs fit within the latency envelopes of practical audits. These results support the view that verifiable fine-tuning can be deployed today in centralized and decentralized pipelines to close the trust gap around LLM adaptation.